# WeatherGFM: Learning A Weather Generalist Foundation Model via In-context Learning

### v1

@@ 4.4 Ablation Studies and Explorations
Scaling law of weather foundation model. To evaluate the impact of data and model scale on performance, we compared single-task models, the base version of our WeatherGFM, and its large version. We established a baseline using a 30M parameter ViT under a single-task with 0.5 million samples. Subsequently, in a multi-task setting with 4 million samples, our model was configured with a base version of 110M and a large version of 330M parameters. Figure 7 illustrates that improvements in performance on various tasks are achieved with the increase of model and data scale. In SR task, we observe that scaling up both the data and the model is essential for performance gains.
Figure 7: The effect of dataset size and model sizes. ViT-ST: single-task ViT trained on 0.5 million samples. Base: our WeatherGFM with 100 M parameters trained on 4 million samples. Large: our WeatherGFM with 330 M parameters trained on 4 million samples.
@@ Appendix A Details of Weather Understanding Tasks
Weather forecasting. Radar echo extrapolation aims to forecast data for the subsequent 1-2 hours utilizing observations from past moments (Gao et al., 2024). This task, similar to precipitation nowcasting, plays a significant role in predicting local weather conditions. It can directly impact traffic plans, disaster warnings, and energy management. Likewise, meteorological satellite image extrapolation is crucial for monitoring and analyzing meteorological conditions. Based on the SEVIR dataset, we consider two weather forecasting tasks: radar echo extrapolation Gong et al. (2024a) and satellite image extrapolation (Shukla et al., 2011). Our weather prediction tasks incorporate observations from the hour before (0, 30, 60, and 90 minutes past) and the hour ahead (120 and 180 minutes into the future) for both radar and satellite IR-069 extrapolation. Consequently, for this task, the SEVIR data was extracted and processed to generate 135,696 sequences for training, along with an independent set of 1,200 sequences to validate/test the fitted model.
@@
Weather super-resolution (SR). Weather spatial super-resolution task (Veillette et al., 2020; Hu et al., 2018; 2019; Li et al., 2019) generates a high-resolution image from a low-resolution(LR) image, while temporal super-resolution predicts a high-resolution(HR) image based on two consecutive observed input images. We take into consideration three weather super-resolution tasks: spatial SR for satellite IR-069, spatial SR for radar VIL, and temporal SR for radar VIL with a one-hour interval. We utilize the SEVIR dataset as the source of the HR image. To obtain the LR image, we employ the “Bicubic” interpolation approach, which is commonly used in vision image SR (Zhang et al., 2019; 2021). In the context of meteorology, this is analogous to statistical downscaling, as described in statistical downscaling (Vandal et al., 2017). Specifically, for the VIL image, given that its original resolution is 384×384, we resize it to 256×256 to serve as the HR image and resize it to 64×64 to function as the LR image, thereby implementing a 4x super-resolution task. For the IR-069 image, since its original image has a resolution of 196×196, we resize it to 256×256 to be the HR image and resize it to 64×64 to be the LR image, thus carrying out a 3x super-resolution task. For each spatial SR for satellite tasks, the SEVIR data was extracted and processed to yield 542,784 images for training, along with an independent set of 4,800 images for validating/testing. For the weather temporal SR, we use the radar VIL image at 1 hour (0 and 60 minutes) as the input to predict the radar VIL image at 30 minutes. For the temporal SR task, the SEVIR data was further extracted and processed to generate 407,088 sequences for training, along with an independent set of 3,600 sequences to validate/test the fitted model.
@@
Weather image translation. Weather image translation involves converting observation data (e.g., satellite data) to a desired weather image (Veillette et al., 2020). For example, depictions of storms obtained from weather radar are extremely important. However, most areas of the world do not have access to ground-based radar. It is useful for generating weather radar images of storm depictions from satellite observation (Veillette et al., 2020). We consider three weather image translation tasks based on SEVIR dataset: translate geostationary IR-069 to geostationary IR-107 data (GEOS-IR2GEOS-IR), geostationary IR-069 to geostationary Visible data (GEOS-IR2GEOS-Vis), translate geostationary IR-069 and IR-107 to radar VIL data (GEOS-IR2Radar). In addition, we add a weather image translation task for environment monitoring: Translate geostationary $NO_{2}$ data to polar-orbiting satellites $NO_{2}$ data (GEOS2POES-$NO_{2}$) based on POMINO-TROPOMI (Liu et al., 2020) and GEOS-CF (Keller et al., 2021) product. For the image translation tasks based on SEVIR dataset, we split SEVIR into 542,784 training samples, 4,800 validation samples and 4,800 test samples. For translating geostationary $NO_{2}$ data, the input images are sourced from GEMS as well as the GEOS-CF datasets, while the output images are obtained from the TROPOMI dataset. Each modality has 20,000 images with a resolution of 256×256. Among them, we allocate 18,000 images as the training set, 1,000 images as the validation set, and 1,000 images as the test set.
@@
Appendix B Effects of Visual Prompts
To assess the effectiveness of visual prompts within the WeatherGFM framework, we conducted a series of tests across various scenarios, including GOES2Radar, Radar Extrapolation, GOES-IR2GOES-IR, Radar Spatial SR, Radar Temporal SR, and Deblur tasks. In our main paper, we detail a thorough process of prompt selection, curating 20 unique prompts for each task and subsequently reporting the most favorable quantitative outcomes. The comprehensive results of these prompt variations on several representative tasks are presented in the tables.
By examining Table 5, we observed that prompts have a significant impact on the GOES2Radar and Radar Extrapolation tasks. Consequently, we conducted further experiments on these two tasks. The results are shown in Table 4. In these experiments, “random prompt” refers to prompts selected randomly, while “high prompt” refers to prompts selected from a high-quality prompt base derived from radar data. Specifically, we grouped 100 events and selected samples that contained values exceeding a threshold of 50 within each group, totaling 113 samples. Prompts were then randomly selected from this high-quality base for the experiments. “Searched prompt” refers to a method where prompts are selected based on the RMSE metric calculated from the input images to find similar prompts.
Tasks GOES2Radar Radar Extrapolation
Appendix C More details of scaling law
Appendix D More Visual Results
To comprehensively assess the performance of WeatherGFM, we present a range of qualitative visual results across various tasks, including weather forecasting, weather super-resolution, and weather image translation. Additionally, we conduct a comparative evaluation against the unified ViT-large model, as well as single-task ViT and UNet models. The visual outputs and comparisons are thoughtfully illustrated in Figure 10.
WeatherGFM’s proficiency in generating visually appealing outputs is readily evident in the presented results. Notably, the visual quality surpasses that of the baseline models. However, the significance of WeatherGFM’s capability extends beyond visual quality. Its distinctive strength lies in its ability to effectively handle a wide array of image enhancement tasks and image detection. This marks a noteworthy distinction from traditional models, which often struggle to concurrently address such a diverse spectrum of tasks.
@@ Appendix D More Visual Results
Appendix E Additional Quantitative Results
In this section, we present more quantitative results on various weather understanding tasks. The results include the Probability of Detection (POD) and False Alarm Rate (FAR) metrics for different models across multiple tasks. Table 6 provides a detailed comparison of these metrics for the UNet, ViT, and WeatherGFM models. AVG.POD and AVG.FAR represent the mean scores across different thresholds for various tasks. For radar output tasks, the thresholds are [16, 74, 133, 160, 181, 219]; for GEOS-visible output tasks, the thresholds are [2000, 3200, 4400, 5600, 6800]; and for GEOS-IR069 output tasks, the thresholds are [-4000, -5000, -6000, -7000].

### v2

@@ 4.4 Ablation Studies and Explorations
Scaling law of weather foundation model. To evaluate the impact of data and model scale on performance, we compared single-task models, the base version of our WeatherGFM, and its large version. We established a baseline using a 30M parameter ViT under a single-task with 0.5 million samples. Subsequently, in a multi-task setting with 4 million samples, our model was configured with a base version of 110M and a large version of 330M parameters. Figure 7 illustrates that improvements in performance on various tasks are achieved with the increase of model and data scale. In specific tasks like radar super-resolution, we observe that scaling up both the data and the model is essential for performance gains.
Figure 7: The effect of model sizes. ViT-ST: single-task ViT trained on 0.5 million samples. Base: our WeatherGFM with 100 M parameters trained on 4 million samples. Large: our WeatherGFM with 330 M parameters trained on 4 million samples.
@@ Appendix A Details of Weather Understanding Tasks
Table 4: Overview of Model Inputs, Outputs, and Prompt Formats
Task Dataset Prompt Format Prompt Input Prompt Output Input Output
Radar Spatial SR Sevir Single Modal VIL LR VIL HR VIL LR VIL HR
Satellite Spatial SR Sevir Single Modal IR-069 LR IR-069 HR IR-069 LR IR-069 HR
Radar Temporal SR Sevir Single Modal VIL (0,60min) VIL 30min VIL (0,60min) VIL 30min
Deblur Sevir Single Modal VIL (Earthformer) VIL (Ground Truth) VIL (Earthformer) VIL (Ground Truth)
GEOS-IR2Radar Sevir Cross Modal IR-069,IR-107 VIL IR-069,IR-107 VIL
GEOS-IR2GEOS-IR Sevir Cross Modal IR-069 IR-107 IR-069 IR-107
GEOS-IR2GEOS-Vis Sevir Cross Modal IR-069 VIS IR-069 VIS
GEOS2POES-NO2 POMINO Cross Modal GEMS, GEOS-CF TROPOMI GEMS, GEOS-CF TROPOMI
Satellite extrapolation Sevir Time-series Modal IR-069 (0,30,60,90min) IR-069 (120,180min) IR-069 (0,30,60,90min) IR-069 (120,180min)
Radar extrapolation Sevir Time-series Modal VIL (0,30,60,90min) VIL (120,180min) VIL (0,30,60,90min) VIL (120,180min)
@@
Weather forecasting. Radar echo extrapolation aims to forecast data for the subsequent 1-2 hours utilizing observations from past moments (Gao et al., 2024). This task, similar to precipitation nowcasting, plays a significant role in predicting local weather conditions. It can directly impact traffic plans, disaster warnings, and energy management. Likewise, meteorological satellite image extrapolation is crucial for monitoring and analyzing meteorological conditions. Based on the SEVIR dataset, we consider two weather forecasting tasks: radar echo extrapolation Gong et al. (2024) and satellite image extrapolation (Shukla et al., 2011). Our weather prediction tasks incorporate observations from the hour before (0, 30, 60, and 90 minutes past) and the hour ahead (120 and 180 minutes into the future) for both radar and satellite IR-069 extrapolation. Consequently, for this task, the SEVIR data was extracted and processed to generate 135,696 sequences for training, along with an independent set of 1,200 sequences to validate/test the fitted model.
@@
Weather super-resolution (SR). Weather spatial super-resolution task (Veillette et al., 2020) generates a high-resolution image from a low-resolution(LR) image, while temporal super-resolution predicts a high-resolution(HR) image based on two consecutive observed input images. We take into consideration three weather super-resolution tasks: spatial SR for satellite IR-069, spatial SR for radar VIL, and temporal SR for radar VIL with a one-hour interval. We utilize the SEVIR dataset as the source of the HR image. To obtain the LR image, we employ the "Bicubic" interpolation approach, which is commonly used in vision image SR. In the context of meteorology, this is analogous to statistical downscaling, as described in statistical downscaling (Vandal et al., 2017). Specifically, for the VIL image, given that its original resolution is 384×384, we resize it to 256×256 to serve as the HR image and resize it to 64×64 to function as the LR image, thereby implementing a 4x super-resolution task. For the IR-069 image, since its original image has a resolution of 196×196, we resize it to 256×256 to be the HR image and resize it to 64×64 to be the LR image, thus carrying out a 3x super-resolution task. For each spatial SR for satellite tasks, the SEVIR data was extracted and processed to yield 542,784 images for training, along with an independent set of 4,800 images for validating/testing. For the weather temporal SR, we use the radar VIL image at 1 hour (0 and 60 minutes) as the input to predict the radar VIL image at 30 minutes. For the temporal SR task, the SEVIR data was further extracted and processed to generate 407,088 sequences for training, along with an independent set of 3,600 sequences to validate/test the fitted model.
@@
Weather image translation. Weather image translation involves converting observation data (e.g., satellite data) to a desired weather image (Veillette et al., 2020). For example, depictions of storms obtained from weather radar are extremely important. However, most areas of the world do not have access to ground-based radar. It is useful for generating weather radar images of storm depictions from satellite observation (Veillette et al., 2020). We consider three weather image translation tasks based on SEVIR dataset: translate geostationary IR-069 to geostationary IR-107 data (GEOS-IR2GEOS-IR), geostationary IR-069 to geostationary Visible data (GEOS-IR2GEOS-Vis), translate geostationary IR-069 and IR-107 to radar VIL data (GEOS-IR2Radar). In addition, we add a weather image translation task for environment monitoring: Translate geostationary $NO_{2}$ data to polar-orbiting satellites $NO_{2}$ data (GEOS2POES-$NO_{2}$) based on POMINO-TROPOMI product (Liu et al., 2020). For the image translation tasks based on SEVIR dataset, we split SEVIR into 542,784 training samples, 4,800 validation samples and 4,800 test samples. For translating geostationary $NO_{2}$ data,
@@
Appendix B Implementation details
B.1 Implementation details and hyperparameters
The hyperparameters for WeatherGFM in our experiments is shown in Table refhyperparameters. The L1 loss is employed as the loss function. For optimization, the AdamW optimizer with a cosine learning rate scheduler is utilized. The base learning rate is 1e-4. The batch size is 20 and the accumulation gradient iterations are 4. We use 16 Nvidia A100 GPUs for training. A total of 50 epochs are executed. We leverage fp16 floating point precision in our model.
Table 5: Default hyperparameters of WeatherGFM
Hyperparameter Meaning Large Base
$p$ Patch size 16 16
Encoder dimension Encoder Embedding dimension $1024$ $768$
Decoder dimension Decoder Embedding dimension $512$ $512$
Encoder depth Number of Encoder blocks $24$ $12$
Decoder depth Number of Encoder blocks $8$ $8$
Encoder Heads Encoder’s attention heads $16$ $12$
Decoder Heads Decoder’s attention heads $16$ $16$
MLP ratio The hidden dimension of the MLP layer in a ViT block $4$ $4$
Masked ratio Percentage of the masked target data $75\%$ $75\%$
B.2 ViT Hyperparameters
We borrow our ViT implementation from (Beyer et al., 2022). We use the following hyperparameters for ViT in all of our experiments.
Table 6: Hyperparameters of ViT
Hyperparameter Meaning Value
$p$ Patch size 16
Dimension Embedding dimension $512$
Depth Number of Encoder blocks $16$
Heads Encoder’s attention heads $8$
MLP dim Encoder’s attention heads $1024$
B.3 UNet Hyperparameters
We use the following hyperparameters for UNet in all of our experiments.
Table 7: Hyperparameters of UNet
Hyperparameter Meaning Value
Padding size Padding size of each convolution layer $1$
Kernel size Kernel size of each convolution layer $3$
Stride Stride of each convolution layer $1$
Channel multiplications Number of output channels for Down and Up blocks [1, 2, 4, 8, 8]
Blocks Number of blocks $3$
Use attention If use attention in Down and Up blocks False
Dropout Dropout rate 00
Inner channel Number of channels in the intermediate layers $64$
Appendix C Extendability
To assess the generalizability of our framework, we also utilize the ERA5 dataset (Hersbach et al., 2020). With the introduction of new meteorological variables, we add a new patch embedding layer for each of these variables to the original model. Following the approach of ClimaX (Nguyen et al., 2023), we use a cross-attention module to aggregate all variable embeddings into a single vector. This allows us to handle the task using the single-modal mode in WeatherGFM. Specifically, we employ an MLP layer to align the embeddings for this task. In line with ClimaX (Nguyen et al., 2023), we use 48 ECMWF (European Centre for Medium-Range Weather Forecasts) variables as input and evaluate the performance of WeatherGFM using the temperature at 2 meters above ground (T2m). We consider seven lead times: 6 hours and 1, 3, 5, 7 days, covering a range from nowcasting to short- and medium-range forecasting. Instead of training separate models for each target variable, our WeatherGFM is trained once to predict all variables across all lead times simultaneously. During fine-tuning, we randomize the lead time from 6 hours to 7 days. Table 8 shows that our approach significantly outperforms ClimaX in the 120-hour and 168-hour forecast results, even surpassing the IFS method of ECMWF. Our method achieves comparable results to ClimaX at other lead times. It’s worth noting that Climax fine-tunes the pre-train model for each lead time, meaning Climax requires N models for weather forecasting at N lead times. In contrast, our universal model can handle all lead time tasks with just a single model, without the need for task-specific fine-tuning. Moreover, the Climax method trained for 100 epochs using 80 V100 GPUs, while our generalist model trained for 20 epochs using 8 A100 GPUs. This indicates that our generalist model converges much faster than Climax.
Table 8: Comparison of RMSE and ACC across different lead times for IFS, ClimaX,and ours WeatherGFM on t2m variable. ClimaX views predicting at each lead time as a separate task and fine-tunes a separate model for every individual task. In contrast, our WeatherGFM utilizes a single model to deal with all of these tasks.
Lead Time RMSE $\downarrow$ ACC $\uparrow$
[hr.] IFS ClimaX WeatherGFM IFS ClimaX WeatherGFM
6 0.97 1.11 1.08 0.99 0.98 0.98
24 1.02 1.19 1.23 0.99 0.97 0.97
72 1.30 1.47 1.56 0.98 0.96 0.96
120 1.71 1.83 1.68 0.96 0.94 0.95
168 2.23 2.17 1.76 0.93 0.91 0.94
@@ Appendix C Extendability
Tasks GOES2Radar Radar Extrapolation
Table 11: The effect of multi-task training. ViT-ST: single-task ViT. WeatherGFM-4Tasks: our WeatherGFM trained on 4 tasks. WeatherGFM-10Tasks: our WeatherGFM trained on full tasks.
Tasks GOES2Radar Radar Temporal SR
Metrics RMSE CSI/74 CSI/133 CSI/181 CSI/219 RMSE CSI/74 CSI/133 CSI/181 CSI/219
ViT-ST 0.445 0.424 0.242 0.134 0.045 0.333 0.585 0.366 0.215 0.063
WeatherGFM-4Tasks 0.460 0.443 0.263 0.166 0.059 0.353 0.576 0.355 0.209 0.074
WeatherGFM-10Tasks 0.436 0.447 0.266 0.157 0.053 0.327 0.597 0.376 0.217 0.073
Tasks Radar Spatial SR GOES2GOES
Metrics RMSE CSI/74 CSI/133 CSI/181 CSI/219 RMSE CSI/-6K CSI/-4K CSI/0 CSI/2K
ViT-ST 0.120 0.820 0.703 0.573 0.387 0.257 0.987 0.972 0.809 0.136
WeatherGFM-4Tasks 0.120 0.831 0.714 0.574 0.380 0.317 0.994 0.968 0.766 0.148
WeatherGFM-10Tasks 0.121 0.831 0.712 0.570 0.375 0.310 0.993 0.968 0.808 0.222
Appendix D Effects of Multi-task Training
In order to assess the influence of multi-task training on performance, we contrasted the versions of our WeatherGFM that were trained on 4 tasks and 10 tasks respectively. Additionally, for the purpose of comparison with WeatherGFM, we also made use of the single-task Vision Transformer (ViT). As shown in the table 11, WeatherGFM-4tasks is trained on four tasks: Radar Temporal SR, GOES2GOES, GOES2Radar, and Radar Spatial SR. It uses more data and encompasses more tasks than the ViT-ST specialized model, which is trained on these four tasks separately. However, WeatherGFM-4tasks utilizes less data and fewer tasks compared to WeatherGFM-10tasks, which is trained on ten tasks. The results indicate that for radar image generation tasks (Radar Temporal SR, GOES2Radar, and Radar Spatial SR), both WeatherGFM-4tasks and WeatherGFM-10tasks outperform ViT-ST. This is likely because most of the selected tasks are related to radar image generation. However, for the satellite image generation task (GOES2GOES), WeatherGFM-4tasks does not perform as well. This suggests that multi-task learning of similar tasks can enhance the model’s performance on those tasks.
Appendix E More details of scaling law
Appendix F Effects of Visual Prompts
To assess the effectiveness of visual prompts within the WeatherGFM framework, we conducted a series of tests across various scenarios, including GOES2Radar, Radar Extrapolation, GOES-IR2GOES-IR, Radar Spatial SR, Radar Temporal SR, and Deblur tasks. In our main paper, we detail a thorough process of prompt selection, curating 20 unique prompts for each task and subsequently reporting the most favorable quantitative outcomes. The comprehensive results of these prompt variations on several representative tasks are presented in the tables. By examining Table 9, we observed that prompts have a significant impact on the GOES2Radar and Radar Extrapolation tasks. Consequently, we conducted further experiments on these two tasks. The results are shown in Table 10. In these experiments, “random prompt” refers to prompts selected randomly, while “high prompt” refers to prompts selected from a high-quality prompt base derived from radar data. Specifically, we grouped 100 events and selected samples that contained values exceeding a threshold of 50 within each group, totaling 113 samples. Prompts were then randomly selected from this high-quality base for the experiments. “Searched prompt” refers to a method where prompts are selected based on the RMSE metric calculated from the input images to find similar prompts.
Appendix G Additional Quantitative Results
In this section, we present more quantitative results on various weather understanding tasks. Table 12 provides quantitative evaluation on OOD tasks. Table 13 provides a detailed comparison of these metrics for the UNet, ViT, and WeatherGFM models. AVG.POD and AVG.FAR represent the mean scores across different thresholds for various tasks. For radar output tasks, the thresholds are [16, 74, 133, 160, 181, 219]; for GEOS-visible output tasks, the thresholds are [2000, 3200, 4400, 5600, 6800]; and for GEOS-IR069 output tasks, the thresholds are [-4000, -5000, -6000, -7000].
Table 12: Quantitative evaluation on OOD tasks. denotes that WeatherGFM has not been trained or fine-tuned for the tasks listed below and conducts generalized inference directly. # indicate that UNet and ViT undergo supervised training on the corresponding training dataset.
Tasks IR107 Satellite extraplotion IR107 2 IR069
Metrics RMSE CSI/-4K CSI/0 CSI/2K RMSE CSI/-4K CSI/-6K
UNet# 0.991 0.695 0.642 0.074 0.942 0.642 0.910
ViT# 0.413 0.899 0.776 0.245 0.212 0.958 0.986
WeatherGFM 0.389 0.903 0.774 0.244 0.340 0.934 0.986
Tasks Temporal SR at 15min
Metrics RMSE CSI/16 CSI/74 CSI/133 CSI/160 CSI/181 CSI/219
UNet# 0.676 0.211 0.627 0.428 0.351 0.262 0.083
ViT# 0.218 0.838 0.761 0.598 0.525 0.445 0.190
WeatherGFM 0.272 0.814 0.703 0.507 0.419 0.336 0.117
@@ Appendix G Additional Quantitative Results
Appendix H More Visual Results
To comprehensively assess the performance of WeatherGFM, we present a range of qualitative visual results across various tasks, including weather forecasting, weather super-resolution, and weather image translation. Additionally, we conduct a comparative evaluation against the unified ViT-large model, as well as single-task ViT and UNet models. The visual outputs and comparisons are thoughtfully illustrated in Figure 10. WeatherGFM’s proficiency in generating visually appealing outputs is readily evident in the presented results. Notably, the visual quality surpasses that of the baseline models. However, the significance of WeatherGFM’s capability extends beyond visual quality. Its distinctive strength lies in its ability to effectively handle a wide array of image enhancement tasks and image detection. This marks a noteworthy distinction from traditional models, which often struggle to concurrently address such a diverse spectrum of tasks.